Sketch the graph of the function, f defined by


f(x)= |x-3| +[x], x∈ [2,4] where [x] denotes the greatest integer function

1
Expert's answer
2022-01-28T13:26:20-0500
"f(x) = \begin{cases}
 3-x+2, &2\leq x<3 \\
 x-3+3, &3\leq x<4 \\
 x-3+4, &x=4
\end{cases}"

"f(x) = \begin{cases}
 5-x, &2\leq x<3 \\
 x, &3\leq x<4 \\
 5, &x=4
\end{cases}"